### v3

@@ 3.1 Primary introduction of ViT model
${Attention(Q,K,V)}=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}+P\right)V$ (1)
where $Q$,$K$,$V$ are the query, key, and value matrices, respectively, representing the input embeddings. $d_{k}$ denotes the dimensionality of $V$, and $P$ means relative position embedding (RPB),
$P_{(x,y),\left(x^{\prime},y^{\prime}\right)}^{h}=Q_{(x,y),:}\cdot K_{\left(x^{\prime},y^{\prime}\right),:}+B_{\left|x-x^{\prime}\right|,\left|y-y^{\prime}\right|}^{h}\cdot$ (2)
where $B^{h}$ represents the translation-invariant attention bias.
@@
Apart from the aforementioned layer, the ViT model also includes a special learnable positional embedding that encodes the position information of each patch in the image sequence. Two common methods of adding position embedding are: 1) applying RPB to the attention map, as in Eq.1; 2) adding absolute position embedding to the patch embeddings. In the vanilla ViT model, preserving positional information is crucial. However, employing position embeddings can be a bit cumbersome, especially when there are changes in the input patch resolution.
Figure 2: Architecture of proposed Tri-Former. SSP block has three main parts, spectral module, spatial module and 3D convolution module, which is why we call it triplet-structured. Spectral module is responsible for collecting information from different spectrum bands. Spatial module is in charge of collecting information from different space locations. 3D convolution layer is added before the two linear layers to enforce 3D structure information and stabilize training process.
3.2 Architecture of Tri-Former
@@
Overall architecture. As is depicted in Figure 2. The overall framework consists of four stages, including one 3D convolution stage and three transformer stages, as shown in the top half of Figure 2. Stage one comprises a spectrum compression module and a 3D convolution module. The subsequent three stages are constructed with the proposed spectral-spatial parallel (SSP) blocks. Each stage begins with a spectral-spatial downsample layer, followed by $n_{i}$ SSP blocks. Following the classical design strategy, the widths of the four stages are progressively doubled from stage to stage. In this work, the width of the first stage is set to 32 as the starting point for subsequent widening. This design choice enables the network to efficiently process and extract relevant information in a computationally efficient manner.
Spectral-spatial parallel block. SSP block first compresses the spectral dimension of input feature. After that, each feature map $f\in{\mathbb{R}}^{c\times s\times w\times h}$ is splitted to sequence and input to the spatial module and spectral module. Specifically, in the spectral module, the spectral query, spectral key, and spectral value can be expressed as:
$Q_{e}\in{{\mathbb{R}}^{R_{e}\times N_{e}\times L_{e}}},\\ K_{e}\in{{\mathbb{R}}^{R_{e}\times N_{e}\times L_{e}}},\\ V_{e}\in{{\mathbb{R}}^{R_{e}\times N_{e}\times L_{e}}},$ (3)
where $R_{e}=h\times w,N_{e}=s,L_{e}=c$. The corresponding spectral attention feature is:
$Atten_{spe}=\operatorname{LN}\left(\frac{\operatorname{softmax}(Q_{e}\cdot K_{e}^{\mathrm{T}})}{\sqrt{d_{e}}}\cdot V_{e}\right),$ (4)
where softmax is conducted along the last dimension, $\cdot$ denotes matrix multiplication and it is conducted along the last two dimension. In the spatial module, the spatial attention feature $Atten_{spa}$ is calculated following the same equation, but the input feature is splitted in different ways. Specifically, the spatial query, spatial key, and spatial value can be represented as:
$Q_{a}\in{{\mathbb{R}}^{R_{a}\times N_{a}\times L_{a}}},\\ K_{a}\in{{\mathbb{R}}^{R_{a}\times N_{a}\times L_{a}}},\\ V_{a}\in{{\mathbb{R}}^{R_{a}\times N_{a}\times L_{a}}},$ (5)
where $R_{a}=s,N_{a}=h\times w,L_{a}=c$. The computing process in SSP block can be formalized as Eq.6 and Eq.7:
$\begin{split}&{f}_{ssp}=f+\operatorname{reshape}(Atten_{spe},\mathbb{R}^{c\times s\times w\times h})+\operatorname{reshape}(Atten_{spa},\mathbb{R}^{c\times s\times w\times h}),\end{split}$ (6)
${{f}_{out}}=\operatorname{MLP}(\operatorname{Conv}({{f}_{ssp}}))+{{f}_{ssp}}$ (7)
where $\operatorname{reshape}(\cdot,\mathbb{R}^{c\times s\times w\times h})$ means reshape the input feature to $\mathbb{R}^{c\times s\times w\times h})$. LN and MLP means linear layer and multilayer perceptron.
@@
Analysis on computational complexity. The proposed spectral-spatial parallel (SSP) block significantly reduces the computation cost. Instead of directly applying self-attention on three dimensions (two spatial dimensions and one spectral dimension), spectral and spatial information are separately processed. The computation complexity of the vanilla ViT is quantic complex regarding to the length of the input sequence. In 3D architecture HSI, this issue is more serious. As shown in Figure 3, for a input feature $f\in{\mathbb{R}}^{c\times s\times w\times h}$, if we directly compute self-attention on three dimensions, the computation complexity is:
@@
Analysis on receptive field. The enhancements on transformer architecture have significant impacts on their receptive field. The receptive field in the proposed SSP block excludes certain irrelevant elements within its scope, consequently strengthening the attention on pixels with close relationships. Using two-dimensional features as an example, we illustrate the receptive fields of SSP block and that of vanilla transformer in Figure 3. From Figure 3(b), it is evident that the receptive field of the vanilla structure is a square region. But as in Figure 3(a), the proposed SSP block considers pixels that share the same row, column, or spatially adjacent positions with the central pixel. The self-attention operation in the ViT architecture achieves information fusion by computing approximate similarities between each pair of pixels. However, when a pixel is relatively distant from the central pixel, it may not effectively contribute to extracting information from the central pixel. Thus, retaining such distant pixels within the receptive field may indeed introduce noise. Therefore, the receptive field of our proposed module retains only the most crucially related pixels, while excluding irrelevant interference pixels, resulting in a more efficient mechanism.
3.3 Attention-Gated Tuning
@@
Attention-Gated Tuning (AGT) is proposed to enable a transformer-based HSI classification method to leverage cross-sensor or cross-modality data to further enhance performance. As illustrated in Figure 4, AGT serves as a bridge, bridging the gaps between target HSIs and images captured by other sensors. The process of AGT is consists of two sequential stages:
Figure 5: Framework of generating pseudo HSIs data from RGB images.
@@
  • Firstly, the basic model is pretrained with homogeneous and heterogeneous hyperspectral data, even RGB data, as illustrated in the blue part of Figure 4. It is worth noting that there is a significant modal disparity between RGB data and hyperspectral data. Therefore, we initially generate pseudo-hyperspectral data from RGB data to mitigate these modality gaps. Specifically, we employ DRCR-Net [^1] https://github.com/jojolee6513/DRCR-net to create the 32-channel pseudo-HSI data from RGB data Li et al. (2022) and then use these pseudo-HSIs to perform pretraining, as depicted in Figure 5. These generated images possess certain spectral characteristics while harnessing the advantages of extensive natural image data. Experimental results in Section 4.8 demonstrate that the AGT can also benefit from RGB data.
@@
  • Secondly, a tiny auxiliary branch is introduced to the basic model before tuning on the target HSI datasets, as presented in the Figure 4. Corresponding to the $n_{i}$ blocks in the basic model, the auxiliary branch consists of $n_{i}$ lightweight SSP blocks. To reduce the overhead associated with the auxiliary branch, we choose smaller patches, shallower architectures, and narrower modules. Through an attention-gated strategy, the proposed AGT can autonomously select useful information from the basic model to guide auxiliary branch learning. Unlike previous fine tuning methods Jia et al. (2022); Liu et al. (2023, 2021); Sung et al. (2022), where the basic model remains frozen during tuning, we employ an asynchronous cold-hot gradient update strategy, in which the basic model is updated slowly.
@@
In terms of design, AGT has three notable aspects that guarantee its effective functioning in our scenario. To elaborate further:
@@
  • Attention-guided learning. In the proposed AGT, features from each stage of the basic model are introduced to guide the auxiliary branch learning. Specifically, as shown in Figure 6, $f_{bas}$ and $f_{end}$ denote the feature from current stage and last stage of basic model, respectively. Shallow features contain an excess of detailed information, whereas deep features encompass more semantic information. AGT combines shallow and deep features learned from pre-training to provide stronger guidance for HSI classification tasks. The gating mechanism is achieved through cross-attention, as shown in Figure 6. This approach allow the auxiliary branch to selectively fuse useful information from the basic model while filtering out conflict information learned during pretraining.
@@
  • Asynchronous cold-hot gradient update strategy. From comprehensive experiments, we find that completely freezing the foundational model as what is done in RGB image processing and NLP fields Jia et al. (2022); Liu et al. (2023) does not work well in HSI classification. We speculate that this phenomenon arises due to the significant disparities among different HSI datasets. Therefore, a more suitable approach involves gradually fine tuning the basic model, allowing it to adjust appropriately and accommodate new data. Therefore, we propose an asynchronous cold-hot gradient update strategy, where basic model is updated slowly (cold) and the auxiliary branch is updated quickly (hot).
@@
Although the design of AGT is inspired by the core idea of ladder side tuning (LST) Jia et al. (2022), their motivations, key problems, architectures and gradient updating are totally different. 1) AGT aims to alleviate limited label sample problems in HSI classification, whereas LTS focuses on fine tuning large language models (LLM); 2) AGT deals with HSIs collected by various sensors or even datasets from different domains. LST is proposed for datasets from the same modality; 3) AGT uses information from the auxiliary branch to guide feature fusion, which reuse the features from the current stage and the end stage; 4) As depicted in Figure 7(b), AGT employ a cold-hot gradient update strategy , while LST always freeze the basic model. The corresponding experiments in Section IV.E demonstrate the effectiveness of proposed AGT.
Figure 7: Architecture comparison on proposed AGT and related work. (a). LST; (b). AGT.
4 Experiments
4.1 Data Description
In order to make a comprehensive comparison, Indian Pines dataset, Pavia University dataset and Houston University dataset are used in the experiments. The sample distribution information is listed in Table 1.

### v4

@@ III-A Primary introduction of ViT model
${Attention(Q,K,V)}=\operatorname{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}+P\right)V$ (1)
where $Q$,$K$,$V$ are the query, key, and value matrices, respectively, representing the input embeddings. $d_{k}$ denotes the dimensionality of feature, and $P$ denotes the relative position embedding term.
@@
Apart from the aforementioned layer, the ViT model also includes a special learnable positional embedding that encodes the position information of each patch in the image sequence. Two common methods of adding position embedding are: 1) applying RPB to the attention map, as in Eq.1; 2) adding absolute position embedding to the patch embeddings. In the vanilla ViT model, preserving positional information is crucial. However, employing position embeddings can be a bit cumbersome, especially when there are changes in the input patch resolution. As shown in Eq.1, in this paper, we adopt the former one. The RPB $P_{(x,y),(x^{\prime},y^{\prime})}^{m}$ is:
$P_{(x,y),(x^{\prime},y^{\prime})}^{m}=B_{|x-x^{\prime}|,|y-y^{\prime}|}^{m},$ (2)
where $B^{m}\in\mathbb{R}^{(2h-1)\times(2w-1)}$ represents the translation-invariant attention bias for the head $m$, with $h,w$ being the spatial dimensions of the feature map.
Figure 2: Architecture of proposed Tri-Former. SSP block has three main parts: spectral module, spatial module and 3D convolution module. Spectral module is responsible for collecting information from different spectrum bands. Spatial module is in charge of collecting information from different space locations. 3D convolution layer is added before the two linear layers to enforce 3D structure information and stabilize training process.
III-B Architecture of Tri-Former
@@
Overall architecture. As is depicted in Fig. 2, following the common setting adopted by ResNet ConvNext, etc. the overall framework consists of four stages. Inspired by FcaFormer [44], LeViT [45], etc, we adopt early convolution architecture. The overall architecture is one 3D convolution stage followed by three transformer stages, as shown in the top half of Fig. 2. Stage one comprises a spectrum compression module and a 3D convolution module. The subsequent three stages are constructed with the proposed spectral-spatial parallel (SSP) blocks. Each stage begins with a spectral-spatial downsample layer, followed by $n_{i}$ SSP blocks. Following the classical design strategy, the widths of the four stages are progressively doubled from stage to stage. In this work, the width of the first stage is set to 32 as the starting point for subsequent widening. This design choice enables the network to efficiently process and extract relevant information in a computationally efficient manner.
Spectral-spatial parallel block. SSP block first compresses the spectral dimension of input feature. After that, each feature map $f\in{\mathbb{R}}^{c\times s\times w\times h}$ is splitted to sequence and input to the spatial module and spectral module.
Specifically, for input feature $f\in\mathbb{R}^{c\times s\times w\times h}$ where $c$=channels, $s$=spectral bands, $(w,h)$=spatial dimensions:
  • Spectral attention splits features along spectral dimension:
    $Q_{e}$ $=f_{W}^{Q}(f)\in\mathbb{R}^{R_{e}\times N_{e}\times L_{e}}$ (3)
    $K_{e}$ $=f_{W}^{K}(f)\in\mathbb{R}^{R_{e}\times N_{e}\times L_{e}}$
    $V_{e}$ $=f_{W}^{V}(f)\in\mathbb{R}^{R_{e}\times N_{e}\times L_{e}}$
    where $R_{e}=w\times h$ (flattened spatial positions), $N_{e}=s$ (spectral bands), $L_{e}=c$ (channel dimension).
  • Spatial attention splits features along spatial dimensions:
    $Q_{a}$ $=f_{W}^{Q}(f)\in\mathbb{R}^{R_{a}\times N_{a}\times L_{a}}$ (4)
    $K_{a}$ $=f_{W}^{K}(f)\in\mathbb{R}^{R_{a}\times N_{a}\times L_{a}}$
    $V_{a}$ $=f_{W}^{V}(f)\in\mathbb{R}^{R_{a}\times N_{a}\times L_{a}}$
    where $R_{a}=s$ (spectral bands), $N_{a}=w\times h$ (spatial positions), $L_{a}=c$ (channel dimension).
The attention computation follows:
$\mathrm{Atten}_{spe}$ $=\mathrm{LayerNorm}\left(\mathrm{softmax}\left(\frac{Q_{e}K_{e}^{\top}}{\sqrt{d_{e}}}\right)V_{e}\right)$ (5)
$\mathrm{Atten}_{spa}$ $=\mathrm{LayerNorm}\left(\mathrm{softmax}\left(\frac{Q_{a}K_{a}^{\top}}{\sqrt{d_{a}}}\right)V_{a}\right)$
where $\mathrm{softmax}$ applies along the second dimension $N$, and $d$ is the scaling factor corresponding to $L$ dimension.
After these attention computations, we reshape $\mathrm{Atten}_{\mathrm{spe}}$ and $\mathrm{Atten}_{\mathrm{spa}}$ back to $\mathbb{R}^{c\times s\times w\times h}$ and combine them with a residual connection:
$f_{\mathrm{ssp}}=f+\mathrm{reshape}\bigl{(}\mathrm{Atten}_{\mathrm{spe}})+\mathrm{reshape}\bigl{(}\mathrm{Atten}_{\mathrm{spa}})$ (6)
To further enhance structural consistency and stabilize training, we apply a 3D convolution followed by an MLP:
$f_{\mathrm{out}}=f_{\mathrm{ssp}}+\mathrm{MLP}\Bigl{(}\mathrm{Conv3D}\!\bigl{(}f_{\mathrm{ssp}}\bigr{)}\Bigr{)}.$ (7)
[…]
@@
Analysis on computational complexity. The proposed spectral-spatial parallel (SSP) block significantly reduces the computation cost. Instead of directly applying self-attention on three dimensions (two spatial dimensions and one spectral dimension), spectral and spatial information are separately processed. The computation complexity of the vanilla ViT is quantic complex regarding to the length of the input sequence. In 3D architecture HSI, this issue is more serious. As shown in Fig. 3, for a input feature $f\in{\mathbb{R}}^{c\times s\times w\times h}$, if we directly compute self-attention on three dimensions, the computation complexity is:
@@
Analysis on receptive field. The enhancements on transformer architecture have significant impacts on their receptive field. The receptive field in the proposed SSP block excludes certain irrelevant elements within its scope, consequently strengthening the attention on pixels with close relationships. Using two-dimensional features as an example, we illustrate the receptive fields of SSP block and that of vanilla transformer in Fig. 3. From Fig. 3(b), it is evident that the receptive field of the vanilla structure is a square region. But as in Fig. 3(a), the proposed SSP block considers pixels that share the same row, column, or spatially adjacent positions with the central pixel. The self-attention operation in the ViT architecture achieves information fusion by computing approximate similarities between each pair of pixels. However, when a pixel is relatively distant from the central pixel, it may not effectively contribute to extracting information from the central pixel. Thus, retaining such distant pixels within the receptive field may indeed introduce noise. Therefore, the receptive field of our proposed module retains only the most crucially related pixels, while excluding irrelevant interference pixels, resulting in a more efficient mechanism.
III-C Attention-Gated Tuning
@@
Attention-Gated Tuning (AGT) is proposed to enable a transformer-based HSI classification method to leverage cross-sensor or cross-modality data to further enhance performance. As illustrated in Fig. 4, AGT serves as a bridge, bridging the gaps between target HSIs and images captured by other sensors. The process of AGT contains two sequential stages:
@@
  • Firstly, the basic model is pretrained with homogeneous and heterogeneous hyperspectral data, even RGB data, as illustrated in the blue part of Fig. 4. It is worth noting that there is a significant modal disparity between RGB data and hyperspectral data. Therefore, we initially generate pseudo-hyperspectral data from RGB data to mitigate these modality gaps. Specifically, we employ DRCR-Net [^1] https://github.com/jojolee6513/DRCR-net to create the 32-channel pseudo-HSI data from RGB data [46] and then use these pseudo-HSIs to perform pretraining. These generated images possess certain spectral characteristics while harnessing the advantages of extensive natural image data. Experimental results in Section IV.I demonstrate that the AGT can also benefit from RGB data.
@@
  • Secondly, a tiny auxiliary branch is introduced to the basic model before tuning on the target HSI datasets, as presented in the Fig. 4. Corresponding to the $n_{i}$ blocks in the basic model, the auxiliary branch consists of $n_{i}$ lightweight SSP blocks. To reduce the overhead associated with the auxiliary branch, we choose smaller patches, shallower architectures, and narrower modules. Through an attention-gated strategy, the proposed AGT can autonomously select useful information from the basic model to guide auxiliary branch learning. Unlike previous fine tuning methods [47, 48, 49, 19], where the basic model remains frozen during tuning, we employ an asynchronous cold-hot gradient update strategy, in which the basic model is updated slowly.
Figure 5: Internal structure of proposed attention gate.
Fig. 5 shows the internal structure of the attention gate. The auxiliary branch feature $f_{\text{aux}}$ is used to generate the query, while the basic model feature $f_{\text{bas}}$ and the final stage feature $f_{\text{end}}$ construct the key and value:
$Q_{\text{aux}}=W_{q}\cdot f_{\text{aux}},$ (10)
$K_{\text{bas}}=W_{k}\cdot f_{\text{bas}},\quad V_{\text{bas}}=W_{v}\cdot f_{\text{bas}},$ (11)
$K_{\text{end}}=W_{k}\cdot f_{\text{end}},\quad V_{\text{end}}=W_{v}\cdot f_{\text{end}},$ (12)
where $W_{q},W_{k},W_{v}$ are learnable projection matrices. The auxiliary branch selectively integrates information from both the basic model and the final stage using two separate cross-attention modules:
$A_{\text{bas}}=\operatorname{softmax}\left(\frac{Q_{\text{aux}}K_{\text{bas}}^{\top}}{\sqrt{d}}\right)V_{\text{bas}},$ (13)
$A_{\text{end}}=\operatorname{softmax}\left(\frac{Q_{\text{aux}}K_{\text{end}}^{\top}}{\sqrt{d}}\right)V_{\text{end}}.$ (14)
The final representation is computed as:
$A_{\text{gate}}=A_{\text{bas}}+A_{\text{end}}+f_{\text{aux}},$ (15)
$f_{\text{out}}=W_{\text{linear}}\cdot(W_{\text{linear}}\cdot f_{\text{bas}}+A_{\text{gate}}).$ (16)
where $W_{\text{linear}}$ is learnable Linear projection matrices.The gating mechanism effectively selects relevant information from both shallow and deep feature representations, ensuring robust knowledge transfer across sensors and modalities.
Algorithm 1 Pseudo-code of Tri-Former with AGT
Input:
1) Source dataset $\mathcal{D}_{S}=\{(x_{i}^{(S)},y_{i}^{(S)})\}$
2) Target dataset $\mathcal{D}_{T}=\{(x_{j}^{(T)},y_{j}^{(T)})\}$
3) Network hyper-parameters (e.g., patch size, channels $c$, spectral bands $s$, etc.)
Output:Trained model parameters $(\theta_{\mathrm{base}},\theta_{\mathrm{aux}})$   Stage A: Pretrain the Base Tri-Former Model
Initialize base model parameters $\theta_{\mathrm{base}}^{(0)}$
Pre-training on mini-batch $\{(x_{i}^{(S)},y_{i}^{(S)})\}\!\subset\!\mathcal{D}_{S}$
1) Forward pass (cf. Eq.(1)-(7))
2) Compute loss: $\mathcal{L}_{S}=\mathrm{CE}\!\bigl{(}\hat{y}_{i}^{(S)},\,y_{i}^{(S)}\bigr{)}$
3) Backward and update base model: $\theta_{\mathrm{base}}\leftarrow\theta_{\mathrm{base}}-\eta\,\nabla_{\theta_{\mathrm{base}}}\,\mathcal{L}_{S}$
  Stage B: Integrate Auxiliary Branch and Perform AGT
Initialize auxiliary branch parameters $\theta_{\mathrm{aux}}^{(0)}$
Tuning on target dataset mini-batch $\{(x_{j}^{(T)},y_{j}^{(T)})\}\!\subset\!\mathcal{D}_{T}$
1) Forward pass (cf. Eq.(10)–(16)):
2) Compute target loss: $\mathcal{L}_{T}=\mathrm{CE}\bigl{(}\hat{y}_{j}^{(T)},\,y_{j}^{(T)})$
3) Asynchronous cold-hot updates:
$\theta_{\mathrm{aux}}\;\leftarrow\;\theta_{\mathrm{aux}}\;-\;\alpha\,\nabla_{\theta_{\mathrm{aux}}}\!\mathcal{L}_{T}\quad(\text{hot: largerLR }\alpha)$
$\theta_{\mathrm{base}}\;\leftarrow\;\theta_{\mathrm{base}}\;-\;\beta\,\nabla_{\theta_{\mathrm{base}}}\!\mathcal{L}_{T}\quad(\text{cold: smaller LR }\beta\ll\alpha)$
Return: $(\theta_{\mathrm{base}},\theta_{\mathrm{aux}})$
@@
In terms of design, AGT has three notable aspects that guarantee its efficiency in our scenario. To elaborate further:
@@
  • Attention-guided learning. In the proposed AGT, features from each stage of the basic model are introduced to guide the auxiliary branch learning. Specifically, as shown in Fig. 5, $f_{bas}$ and $f_{end}$ denote the feature from current stage and last stage of basic model, respectively. Shallow features contain an excess of detailed information, whereas deep features encompass more semantic information. AGT combines shallow and deep features learned from pre-training to provide stronger guidance for HSI classification tasks. The gating mechanism is achieved through cross-attention, as shown in Fig. 5. This approach allow the auxiliary branch to selectively fuse useful information from the basic model while filtering out conflict information learned during pretraining.
@@
  • Asynchronous cold-hot gradient update strategy. From comprehensive experiments, we find that completely freezing the foundational model as what is done in RGB image processing and NLP fields [47, 48] does not work well in HSI classification. We speculate that this phenomenon arises due to the significant disparities among different HSI datasets. Therefore, a more suitable approach involves gradually fine tuning the basic model, allowing it to adjust appropriately and accommodate new data. Therefore, we propose an asynchronous cold-hot gradient update strategy, where basic model is updated slowly (cold) and the auxiliary branch is updated quickly (hot).
@@
Although the design of AGT is inspired by the core idea of ladder side tuning (LST) [47], their motivations, key problems, architectures and gradient updating are totally different. 1) AGT aims to alleviate limited label sample problems in HSI classification, whereas LTS focuses on fine tuning large language models (LLM); 2) AGT deals with HSIs collected by various sensors or even datasets from different domains. LST is proposed for datasets from the same modality; 3) AGT uses information from the auxiliary branch to guide feature fusion, which reuse the features from the current stage and the end stage; 4) As depicted in Fig. 6(b), AGT employ a cold-hot gradient update strategy , while LST always freeze the basic model. The corresponding experiments in Section IV.G demonstrate the effectiveness of proposed AGT.
III-D Theoretical Analysis from Domain Adaptation Perspective
In this section, we provide a theoretical perspective on Attention-Gated Tuning (AGT) from the perspective of domain adaptation. Cross-sensor or cross-modality hyperspectral data often exhibit significant discrepancies from the data used to pre-train the backbone model. Consequently, naive fine-tuning (either fully or partially) on the new target dataset may lead to suboptimal performance. Below, we show how AGT, through its gating mechanism and asynchronous cold-hot gradient updates, severed as an effective strategy for mitigating domain gaps while retaining the benefits of large-scale pre-trained data.
III-D1 Revisiting the Domain Adaptation Bound
Let $\mathcal{D}_{S}$ denote the source distribution (from which large-scale pretraining data are sampled) and $\mathcal{D}_{T}$ denote the target distribution (the new hyperspectral dataset). In classical domain adaptation theory [50], for any hypothesis $h$ in a hypothesis space $\mathcal{H}$, the error on the target distribution, $\epsilon_{T}(h)$, can be upper-bounded as:
$\epsilon_{T}(h)\;\leq\;\epsilon_{S}(h)\;+\;\frac{1}{2}\,d_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{D}_{S},\mathcal{D}_{T})\;+\;\lambda,$ (17)
where $\epsilon_{S}(h)$ is the error on the source distribution, $d_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{D}_{S},\mathcal{D}_{T})$ measures the divergence between source and target distributions under $\mathcal{H}$, and $\lambda$ reflects the best possible joint error across both domains.
In cross-sensor or cross-modality scenarios, the discrepancy $d_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{D}_{S},\mathcal{D}_{T})$ can be large. Standard fine-tuning faces two major risks:
  • Over-updating: the pre-trained model with large learning rates, thereby forgetting valuable source-domain knowledge and increasing $\epsilon_{S}(h)$,
  • Under-updating: the model with too small learning rate or frozen, failing to adapt to the target domain and keeping $\epsilon_{T}(h)$ suboptimal.
Under this basis, we explain how the proposed AGT design alleviates these issues.
III-D2 AGT as an Information Selection Mechanism
AGT introduces a tiny auxiliary branch in parallel to the base model. We denote: $\theta_{\text{base}}$ as parameters of the pre-trained model, $\theta_{\text{aux}}$ as parameters of the lightweight branch, and let $G(\cdot)$ be the cross-attention-based gating module. Thus, the final hypothesis $h$ can be expressed as:
$h(\mathbf{x};\theta_{\text{base}},\theta_{\text{aux}})\;=\;G\bigl{(}f_{\text{base}}(\mathbf{x};\theta_{\text{base}}),\;f_{\text{aux}}(\mathbf{x};\theta_{\text{aux}})\bigr{)},$ (18)
where $f_{\text{base}}(\mathbf{x};\theta_{\text{base}})$ is the feature map from the base model, and $f_{\text{aux}}(\mathbf{x};\theta_{\text{aux}})$ is the feature map from the auxiliary branch.
  • Selective Retention of Pre-trained Knowledge. Within the gating module, both shallow and deep features from the base model serve as key-value pairs, while the auxiliary branch feature forms the query. Through cross-attention, the auxiliary branch “selects” only the most relevant information from the base model. Irrelevant or conflicting channels are less likely to be activated. Hence, the pre-trained representations learned from large-scale data remain available and are not overwritten, aligning with the core goal of domain adaptation: preserving source knowledge to avoid an excessive increase in $\epsilon_{S}(h)$.
  • Auxiliary Branch for Target-specific Adaptation. For target data $\mathcal{D}_{T}$ that differ significantly from $\mathcal{D}_{S}$, the auxiliary branch $\theta_{\text{aux}}$ can freely learn new spectral-spatial patterns, not fully captured by the base model. By integrating these newly learned features through the gating mechanism in Eq.(18), the overall hypothesis space effectively expands from $\mathcal{H}$ (original base model) to ~=$\widetilde{\mathcal{H}}=\{G(f_{\text{base}},f_{\text{aux}})\}$. This extra flexibility can bridge large domain gaps $d_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{D}_{S},\mathcal{D}_{T})$. From the view of Eq.(17), AGT thus aims to (i) keep $\epsilon_{S}(h)$ low by preserving core parameters and (ii) reduce $d_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{D}_{S},\mathcal{D}_{T})$ with new trainable components $\theta_{\text{aux}}$.
Figure 6: Comparison on proposed AGT and related work. (a). LST; (b). AGT.
III-D3 Asynchronous Cold-Hot Updates for Balanced Optimization
While the above perspective highlights the importance of preserving base-model knowledge, it is equally crucial to adapt efficiently to $\mathcal{D}_{T}$. AGT uses an asynchronous cold-hot gradient update to address this:
$\theta_{\text{aux}}^{(t+1)}$ $=\theta_{\text{aux}}^{(t)}-\alpha\,\nabla_{\theta_{\text{aux}}}\mathcal{L}_{T}\bigl{(}\theta_{\text{aux}}^{(t)},\theta_{\text{base}}^{(t)}\bigr{)},$ (19)
$\theta_{\text{base}}^{(t+1)}$ $=\theta_{\text{base}}^{(t)}-\beta\,\nabla_{\theta_{\text{base}}}\mathcal{L}_{T}\bigl{(}\theta_{\text{aux}}^{(t)},\theta_{\text{base}}^{(t)}\bigr{)},$
where $\alpha\gg\beta$. Consequently,
  • $\theta_{\text{aux}}$ (hot) rapidly adapts to domain-specific cues in $\mathcal{D}_{T}$,
  • $\theta_{\text{base}}$ (cold) is updated more conservatively, preventing catastrophic forgetting of the source knowledge.
Hence, this setup finds a middle ground between the extremes of (i) freezing $\theta_{\text{base}}$ (potentially insufficient for large domain shifts), and (ii) fully fine-tuning with large learning rates (risking an overcorrection that erases pre-trained representations).
III-D4 Summary of Domain Adaptation Insights
From a domain adaptation standpoint, the proposed AGT design can be summarized as follows:
  • Gate-based feature fusion (Eq. (18)) selectively extracts relevant knowledge from the pre-trained base model without inheriting spurious or conflicting information.
  • Auxiliary branch $\theta_{\text{aux}}$ captures new spectral-spatial patterns crucial for the target hyperspectral domain, effectively reducing the distribution gap $d_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{D}_{S},\mathcal{D}_{T})$.
  • Asynchronous cold-hot updates maintain a stable optimization path that balances preserving source knowledge and adapting to the new domain.
As validated by our experiments in Section IV, these ideas collectively enable robust and efficient adaptation, even under challenging cross-sensor or cross-modality conditions where domain discrepancies are pronounced.
IV Experiments
IV-A Data Description
In order to comprehensively evaluate the performance of our proposed method under diverse conditions, we selected six representative datasets. These datasets vary in sensor type, spectral band count, spatial resolution, and scene characteristics. This selection addresses three key challenges: (1) scene specificity (differing land-cover types and background complexities), (2) spatial resolution differences (mixed pixels in low resolution versus fine spatial details in high resolution), and (3) the challenge of effectively fusing spectral and spatial information under varying conditions. Overall, the diversity of the chosen datasets ensures a fair and rigorous evaluation of the proposed method across different imaging scenarios. The sample distribution information is listed in Table  I and Table II, and the detail introduction of these datasets are as follows: